Problem 9-11 Calculating Project Cash Flow from Assets [LO 2]

Cochrane, Inc., is considering a new three-year expansion project that requires an initial fixed asset investment of $1,860,000. The fixed asset will be depreciated straight-line to zero over its three-year tax life, after which time it will be worthless. The project is estimated to generate $1,950,000 in annual sales, with costs of $1,060,000. The project requires an initial investment in net working capital of $150,000, and the fixed asset will have a market value of $175,000 at the end of the project. Assume that the tax rate is 35 percent and the required return on the project is 14 percent.

Requirement 1:

What are the net cash flows of the project for the following years? (Do not round intermediate calculations. Negative amounts should be indicated by a minus sign. Enter your answers in dollars, not millions of dollars (e.g., 1,234,567).)

Year Cash Flow
0 $
1
2
3
Requirement 2:

What is the NPV of the project? (Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars (e.g., 1,234,567). Round your answer to 2 decimal places (e.g., 32.16).)

  NPV $
Answer #1
Year 0 Year 1 Year 2 Year 3
Annual Sales                   -            1,950,000          1,950,000          1,950,000
Less: Costs                   -            1,060,000          1,060,000          1,060,000
Earnings before depreciation                   -                890,000              890,000              890,000
Less: Depreciation                   -                620,000              620,000              620,000
Earnings before tax                   -                270,000              270,000              270,000
Tax at 35%                   -                  94,500                94,500                94,500
Earnings after tax                   -                175,500              175,500              175,500
Add back depreciation                   -                620,000              620,000              620,000
Cash flow from operations                   -                795,500              795,500              795,500
Initial Investment -1860000                         -                           -                           -  
Investment in Net working capital -150000                         -                           -                150,000
After tax cash flow from sale of asset 0                         -                           -                113,750
Net Cash flow -2010000              795,500              795,500          1,059,250
Discount factor at 14% 1 0.877192982 0.769467528 0.674971516
Discounted cash flow -2010000        697,807.02        612,111.42        714,963.58
NPV =    14,882.01

Formulas used in the table above:

\textup{Straight line depreciation} = \frac{\textup{Cost of Asset - Salvage value} }{\textup{Useful life}}

\textup{After tax cash flow from sale of asset} = \textup{Market price} \times (1-\textup{tax rate})

NPV =\textup{ Discounted Cash Inflow - Discounted cash outflow}

\textup{Discount factor} = \frac{1}{\left (1+i \right )^n}
Where,

i = required rate of return
n = number of years.

Other notes:

Depreciation is added back because it is a non-cash expense, but tax deductible.
Investment in Net Working Capital will be released at the end of any project.

Excel formulas:

A B C D Year 0 0 0 0 0 Year 1 1950000 1060000 =C2-C3 =1860000/3 =C4-C5 =C6*35% =C6-C7 =C5 =C8+C9 Year 2 1950000 1060000 =D2-D

If you need any more details or have any doubts, ask me in the comments.